If a soft-drink bottle whose volume is 1.10 L is completely filled with water and then frozen to -10 ∘C, what volume does the ice occupy? Water has a density of 0.997 g/cm3 at 25 ∘C; ice has a density of 0.917 g/cm3 at -10 ∘C.
Answer #1

The density of ice is slightly less than the density of water; hence when water freezes to ice, the volume will be very close to the initial volume of water. Water will fill the soft drink bottle completely, so the volume occupied by the water = 1.10 L = (1.10 L)*(1000 mL/1 L) = 1100 mL.

The density of water is 0.997 gm/cm3 = 0.997 gm/mL at 25°C; the weight of water in the soft drink bottle = (1100 mL)*(0.997 gm/mL) = 1096.7 gm.

The weight of water frozen will remain unchanged, hence the weight of ice = 1096.7 gm.

However, the density of water is 0.917 gm/cm3 = 1096.7 gm/mL at -10°C; therefore, the volume occupied by ice at -10°C = (1096.7 gm)/(0.917 gm/mL) = 1195.9 mL = (1195.9 mL)*(1 L/1000 mL) = 1.195 L ≈ 1.20 L (ans).